A wire has a resistance of 10.0 ohms at 20.00c and 13.1 ohms at 1000c. Obtain a value for its temperature coefficient of resistance.
1
Expert's answer
2018-02-26T09:27:07-0500
The temperature dependence of a wire resistance is given by equation
R(t)=R_0 (1+αt)
So
R(t_1 )=R_0 (1+αt_1 )
R(t_2 )=R_0 (1+αt_2 )
These equations give
α=(R(t_2 )-R(t_1 ))/(t_2 R(t_1 )-t_1 R(t_2 ) )
α=(13.1-10)/(100×10-20×13.1)=0.0042 1/℃
Answer: 0.0042 1/℃